# Learning A Physical-aware Diffusion Model Based on Transformer for Underwater Image Enhancement

### v1

@@ Abstract
Underwater visuals undergo various complex degradations, inevitably influencing the efficiency of underwater vision tasks. Recently, diffusion models were employed to underwater image enhancement (UIE) tasks, and gained SOTA performance. However, these methods fail to consider the physical properties and underwater imaging mechanisms in the diffusion process, limiting information completion capacity of diffusion models. In this paper, we introduce a novel UIE framework, named PA-Diff, designed to exploiting the knowledge of physics to guide the diffusion process. PA-Diff consists of Physics Prior Generation (PPG) Branch and Physics-aware Diffusion Transformer (PDT) Branch. Our designed PPG branch is a plug-and-play network to produce the physics prior, which can be integrated into any deep framework. With utilizing the physics prior knowledge to guide the diffusion process, PDT branch can obtain underwater-aware ability and model the complex distribution in real-world underwater scenes. Extensive experiments prove that our method achieves best performance on UIE tasks.
Index Terms—  Underwater image enhancement, physics model, diffusion model, transformer
1 Introduction
@@
The restoration of images beneath the water’s surface is a pragmatic yet intricate technology within the realm of underwater vision, extensively applied in endeavors such as underwater robotics[1] and tracking underwater objects[2]. Owing to the phenomena of light refraction, absorption, and scattering in underwater surroundings, images captured underwater are often subject to substantial distortion, characterized by diminished contrast and inherent blurriness [3]. Consequently, the acquisition of clear images in underwater scenarios assumes paramount importance, especially in disciplines necessitating interactions with the submerged environment. The primary objective of underwater image enhancement (UIE) lies in the attainment of superior-quality images, achieved through the elimination of scattering effects and rectification of color distortions prevalent in degraded images. UIE stands as an indispensable component for tasks related to vision in the underwater domain.
@@
To address this problem, conventional UIE approaches, which hinge on the intrinsic characteristics of underwater images, have been presented [4, 5, 6]. These techniques delve into the physical aspects contributing to degradation, such as color cast or scattering, seeking to rectify and enhance the underwater images. Nevertheless, these models grounded in physics exhibit restricted representational capacities, rendering them insufficient for comprehensively addressing the intricate physical and optical elements inherent in underwater scenes. Consequently, their efficacy diminishes, yielding suboptimal enhancement outcomes, particularly in the face of highly intricate and diverse underwater scenarios. In recent times, many learning-based methodologies [7, 8, 9, 10] have been introduced to yield superior results. Leveraging the potent feature representation and nonlinear mapping capabilities of neural networks, these methods excel. They can effectively learn the transformation from a degraded image to a clear one through extensive paired training data.
@@
Recently, there has been a surge of interest in image synthesis [11, 12] and restoration tasks [13, 14, 15], with a notable focus on diffusion-based techniques like DDPM [16] and DDIM [17]. This heightened attention can be attributed to the remarkable generative capabilities of diffusion models. These methods can gradually recover images from noisy images generated in the forward diffusion process and achieve high-quality mapping from randomly sampled Gaussian noise to target images in the reverse diffusion process [16], which poses a new perspective for underwater image enhancement task. Tang et al. [7] present an image enhancement approach with diffusion model in underwater scenes, and WF-Diff [18] proposed a underwater image enhancement framework based on frequency domain information and diffusion models. However, these methods fail to consider the physical properties and underwater imaging mechanisms in the diffusion process, limiting their information completion capacity. Moreover, the diffusion models lack awareness regarding the difficulty of enhancing different image regions, a critical consideration for modeling the complex distribution in real-world underwater scenes.
@@
In this paper, we develop a novel physics-aware diffusion model, fully exploiting physical information to guide the diffusion process, called PA-Diff. It mainly consists of two branches: Physics Prior Generation (PPG) Branch and Physics-aware Diffusion Transformer (PDT) Branch. The first branch aims to generate the transmission map and the global background light as physics prior information by utilizing the modified koschmieder light scanning model. The transmission map is exploited as the confidence guidance for the PDT branch which enables our PA-Diff with underwater-aware ability. The PDT branch employs the strong generation ability of diffusion models to restore underwater images, with the guidance of physical information. Specifically, we design a physics-aware diffusion transformer block which contain a physics-aware self-attention (PA-SA) and multi-scale dynamic feed-forward network (MS-FFN) to exploit physics prior information and capture the long-range diffusion dependencies. Our PA-Diff achieves unprecedented perceptual performance in UIE task. Extensive experiments demonstrate that our developed PA-Diff performs the superiority against previous UIE approaches, and ablation study can demonstrate the effectiveness of all contributions.
@@
  • We propose a novel UIE framework based on physics-aware diffusion model, named PA-Diff, which consists of Physics Prior Generation (PPG) Branch and Physics-aware Diffusion Transformer (PDT) Branch. To the best of our knowledge, it is the first the Diffusion model with the guidance of physical knowledge in underwater image enhancement tasks.
  • We design a physics-aware diffusion transformer block, which not only enables PA-Diff with underwater-aware ability to guide the diffusion process, but also captures the long-range diffusion dependencies.
@@
Fig. 1: Overall framework of our proposed PA-Diff. PA-Diff mainly consists of two cooperative branches: Physics Prior Generation (PPG) Branch and Physics-aware Diffusion Transformer (PDT) Branch. Our designed PPG branch is a plug-and-play universal module to produce the prior knowledge of physics, which can be integrated into any deep learning framework. With utilizing the physics prior knowledge to guide the diffusion process, PDT branch can obtain underwater-aware ability and model the complex distribution in real-world underwater scenes.
2 Related Works
2.1 Underwater Image Enhancement
@@
Currently, existing UID methods can be briefly categorized into the physical and deep model-based approaches [4, 6, 7, 8, 9]. Most UID methods based on the physical model utilize prior knowledge to establish models, such as water dark channel priors [4], attenuation curve priors [19], fuzzy priors [20]. In addition, Akkaynak and Treibitz [21] proposed a method based on the revised physical imaging model. The manually established priors restrain the model’s robustness and scalability under the complicated and varied circumstances. Recently, deep learning-based methods [7, 8, 9] achieve acceptable performance, and some complex frameworks are proposed and achieve the-state-of-the-art performance [4, 5, 22]. The previous methods neglect the underwater imaging mechanism and rely only on the representation ability of deep networks. Ucolor [23] combined the underwater physical imaging model in the raw space and designed a medium transmission guided model. ATDCnet [24] attempted to use transmission maps to guide deep neural networks. Therefore, how to fully exploit the knowledge of physics in deep neural networks is a very crucial problem.
2.2 Diffusion Model
@@
Diffusion Probabilistic Models (DPMs) [16, 17] have been widely adopted for conditional image generation [13, 14, 25, 26, 27]. Saharia et al. [28] proposes Palette, which has demonstrated the excellent performance of diffusion models in the field of conditional image generation, including colorization, in-painting and JPEG restoration. Currently, diffusion models was employed to a variety of low-level visual tasks, and gained SOTA performance such as image restoration [29] and low light image enhancement [30]. Recently, Tang et al. [7] presented an image enhancement approach with diffusion model in underwater scenes. WF-Diff [18] proposed a underwater image enhancement framework based on frequency domain information and diffusion models. However, these methods fail to consider the physical properties and underwater imaging mechanisms in the diffusion process, limiting information completion capacity of diffusion models.
3 Methodology
3.1 Overall Framework
@@
The overall framework of PA-Diff is shown in Fig. 1. PA-Diff is designed to leverage the physical properties of underwater imaging mechanisms to guide the diffusion process, which mainly consists of Physics Prior Generation (PPG) Branch and Physics-aware Diffusion Transformer (PDT) Branch. The PPG branch aims to generate the transmission map and the global background light as physics prior information by utilizing the modified koschmieder light scanning model. The transmission map is exploited as the confidence guidance for the PDT branch which enables our PA-Diff with underwater-aware ability. The PDT branch employs the strong generation ability of diffusion models to restore underwater images, with the guidance of physical information. We adopt the diffusion process proposed in DDPM [16] to construct the PDT branch, which can be described as a forward diffusion process and a reverse diffusion process.
@@
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I),$ (1)
@@
$(x_{t}=\sqrt{\alpha_{t}}x_{t-1}+\sqrt{1-\alpha_{t}}\epsilon_{t-1},\quad\epsilon_{t-1}\sim\mathcal{N}(0,\mathcal{Z}).\text{(}$2)
@@
$q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha_{t}}}x_{0},(1-\bar{\alpha_{t}})I).$ (3)
@@
$p_{\theta}(x_{t-1}|x_{t},x_{c})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},x_{c},t),\sigma_{t}^{2}\mathcal{Z}),$ (4)
where $x_{c}$ refers to the conditional image ${I}$ (underwater input image). $\boldsymbol{\mu}_{\theta}(x_{t},x_{c},t)$ and $\sigma_{t}^{2}$ are the mean and variance from the estimate of step t, respectively. We follow the setup of [16], they can be expressed as:
$\mu_{\theta}(x_{t},x_{c},t)=\frac{1}{\sqrt{\alpha}_{t}}(x_{t}-\frac{\beta_{t}}{(1-\overline{\alpha}_{t})}\epsilon_{\theta}(x_{t},x_{c},t)),$ (5)
$\sigma_{t}^{2}=\frac{1-\overline{\alpha}_{t-1}}{1-\overline{\alpha}_{t}}\beta_{t},$ (6)
@@
$L_{dm}(\theta)=\|\epsilon-\epsilon_{\theta}(\sqrt{\overline{\alpha}_{t}}x_{0}+\sqrt{1-\overline{\alpha}_{t}}\epsilon,x_{c},t)\|.$ (7)
3.2 Physics Prior Generation
Mathematically, the underwater images can be expressed by the modified Koschmieder light scanning model:
$I^{c}(x)=J^{c}(x)T^{c}(x)+(1-T^{c}(x))B^{c},$ (8)
where $c\in\{R,G,B\}$ is the color channel, $I^{c}$ is the observed underwater image, $J^{c}$ is the restored clean image, $T^{c}(x)$ is the medium transmission map, and $B^{c}$ is the global background scattered light. In Physics Prior Generation (PPG) branch, our goal is to produce $T^{c}(x)$ and $B^{c}$ from the underwater image $I^{c}$, exploiting the physics model.
Network architecture of PPG branch is shown in Fig. 1, and consists of two sub-networks: transmission map generation sub-network $\mathcal{T}(x)$ and background light generation sub-network $\mathcal{B}(x)$. The two sub-networks consist of duplicated Dynamic Convolutions (DC) [31]. DC dynamically combines numerous parallel convolutional kernels according to their attention, thereby enhancing the model’s complexity and performance without augmenting the network’s depth or breadth.

### v2

@@ Abstract.
Underwater visuals undergo various complex degradations, inevitably influencing the efficiency of underwater vision tasks. Recently, diffusion models were employed to underwater image enhancement (UIE) tasks, and gained SOTA performance. However, these methods fail to consider the physical properties and underwater imaging mechanisms in the diffusion process, limiting information completion capacity of diffusion models. In this paper, we introduce a novel UIE framework, named PA-Diff, designed to exploiting the knowledge of physics to guide the diffusion process. PA-Diff consists of Physics Prior Generation (PPG) Branch, Implicit Neural Reconstruction (INR) Branch, and Physics-aware Diffusion Transformer (PDT) Branch. Our designed PPG branch aims to produce the prior knowledge of physics. With utilizing the physics prior knowledge to guide the diffusion process, PDT branch can obtain underwater-aware ability and model the complex distribution in real-world underwater scenes. INR Branch can learn robust feature representations from diverse underwater image via implicit neural representation, which reduces the difficulty of restoration for PDT branch. Extensive experiments prove that our method achieves best performance on UIE tasks. The code is available at https://github.com/chenydong/PA-Diff.
Underwater image enhancement, Physics model, Diffusion model, Transformer
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologiesccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Computer vision
1. Introduction
@@
The restoration of images beneath the water’s surface is a pragmatic yet intricate technology within the realm of underwater vision, extensively applied in endeavors such as underwater robotics(McMahon and Plaku, 2021) and tracking underwater objects(de Langis and Sattar, 2020). Owing to the phenomena of light refraction, absorption, and scattering in underwater surroundings, images captured underwater are often subject to substantial distortion, characterized by diminished contrast and inherent blurriness (Akkaynak et al., 2017). Consequently, the primary objective of underwater image enhancement (UIE) lies in the attainment of superior-quality images, achieved through the elimination of scattering effects and rectification of color distortions prevalent in degraded images.
@@
To address this problem, conventional UIE approaches, which hinge on the intrinsic characteristics of underwater images, have been presented (Peng and Cosman, 2017; Jr. et al., 2013; Peng et al., 2018). These techniques delve into the physical aspects contributing to degradation, such as color cast or scattering, seeking to rectify and enhance the underwater images. Nevertheless, these models grounded in physics exhibit restricted representational capacities, rendering them insufficient for comprehensively addressing the intricate physical and optical elements inherent in underwater scenes. Consequently, their efficacy diminishes, yielding suboptimal enhancement outcomes, particularly in the face of highly intricate and diverse underwater scenarios.
In recent times, many learning-based methodologies (Tang et al., 2023; Peng et al., 2023; Li et al., 2020; Fabbri et al., 2018) have been introduced to yield superior results. Leveraging the potent feature representation and nonlinear mapping capabilities of neural networks, these methods excel. Deep learning-based methods can be roughly categorized into two classes: physics-free end-to-end methods (Jiang et al., 2022; Peng et al., 2023) and physics-aware methods (Li et al., 2021; Mu et al., 2023b; Fu et al., 2022a). Regarding the physics-free end-to-end methods, most of them usually use ground-truth images as labels to enforce L1/L2 distance-based consistency and also introduce various regularizations as additional constraints to deal with the ill-posed property. Although these methods achieve decent performance, they mainly consider the model in the raw image space, which does not fully explore the useful physics-aware feature information for UIE. Regarding the physics-aware deep methods, most of them utilize the atmospheric scattering model. Existing UIE approaches (Li et al., 2021; Fu et al., 2022a) develop deep neural networks to estimate the transmission and the background scattered light, and then follow the traditional UIE methods to restore underwater images. However, these approaches do not correct the errors of the atmospheric light estimated by neural networks, which seriously affects the quality of the images (Ye et al., 2022). Therefore, there is a crucial problem to adequately exploit the physical information and effectively incorporate them into a unified UIE deep framework.
@@
Recently, there has been a surge of interest in image synthesis (Rombach et al., 2022; Saharia et al., 2022b; Lu et al., 2024, 2023; Zhou et al., 2024) and restoration tasks (Choi et al., 2021; Wang et al., 2023; Yi et al., 2023; Zhou et al., 2023), with a notable focus on diffusion-based techniques like DDPM (Ho et al., 2020) and DDIM (Song et al., 2021). This heightened attention can be attributed to the remarkable generative capabilities of diffusion models.These methods can gradually recover images from noisy images generated in the forward diffusion process and achieve high-quality mapping from randomly sampled Gaussian noise to target images in the reverse diffusion process (Ho et al., 2020), which poses a new perspective for underwater image enhancement task. Tang et al. (Tang et al., 2023) present an image enhancement approach with diffusion model in underwater scenes, and WF-Diff (Zhao et al., 2024) proposed a underwater image enhancement framework based on frequency domain information and diffusion models. UIEDP propose a UIE framework with Diffusion Prior, treating UIE as a posterior distribution sampling process of clear images conditioned on degraded underwater inputs(Du et al., 2023). However, these methods fail to consider the physical properties and underwater imaging mechanisms in the diffusion process, limiting their information completion capacity for the diffusion model. Moreover, the diffusion models lack awareness regarding the difficulty of enhancing different image regions, a critical consideration for modeling the complex distribution in real-world underwater scenes.
@@
In this paper, we develop a novel physics-aware diffusion model, fully exploiting physical information to guide the diffusion process, called PA-Diff. It mainly consists of three branches: Physics Prior Generation (PPG) Branch, Implicit Neural Reconstruction (INR) Branch, and Physics-aware Diffusion Transformer (PDT) Branch. The PPG branch aims to generate the transmission map and the global background light as physics prior information by utilizing the modified koschmieder light scanning model (Akkaynak and Treibitz, 2019a). The transmission map is exploited as the confidence guidance for the PDT branch which enables our PA-Diff with underwater-aware ability. The global background light can be regarded as a prior condition on both the type and degree of degradation for the PDT Branch. Inspired by the latest works (Chen et al., 2024; Yang et al., 2023a) based on INR that can encode an image as a continuous function and learn a more robust degradation representation, we further incorporate the INR into our proposed framework to learn robust feature representations from diverse underwater image. The PDT branch employs the strong generation ability of diffusion models to restore underwater images, with the guidance of physical information. Specifically, we design a physics-aware diffusion transformer block which contain a physics-aware self-attention (PA-SA), physics perception unit (PPU) and gated multi-scale feed-forward network (GM-FFN) to exploit physics prior information and capture the long-range diffusion dependencies. Our PA-Diff achieves unprecedented perceptual performance in UIE task. Extensive experiments demonstrate that our developed PA-Diff performs the superiority against previous UIE approaches, and ablation study can demonstrate the effectiveness of all contributions.
@@
  • We propose a novel UIE framework based on physics-aware diffusion model, named PA-Diff, which consists of Physics Prior Generation (PPG) Branch, Implicit Neural Reconstruction (INR) Branch, and Physics-aware Diffusion Transformer (PDT) Branch. To the best of our knowledge, it is the first the Diffusion model with the guidance of physical knowledge in underwater image enhancement tasks.
  • We incorporate Implicit Neural Reconstruction into our proposed PA-Diff to learn robust feature representations from diverse underwater image, which reduces the difficulty of restoration for diffusion models.
  • We design a physics-aware diffusion transformer block, which not only enables PA-Diff with underwater-aware ability to guide the diffusion process, but also captures the long-range diffusion dependencies. Moreover, we also propose a physics perception unit (PPU), which can explore the beneficial feature-level physics information for underwater image enhancement.
@@
Figure 1. Overall framework of our proposed PA-Diff. PA-Diff mainly consists of three cooperative branches: Physics Prior Generation (PPG) Branch, Implicit Neural Reconstruction (INR) Branch, and Physics-aware Diffusion Transformer (PDT) Branch. Our designed PPG branch aims to produce the prior knowledge of physics. With utilizing the physics prior knowledge to guide the diffusion process, PDT branch can obtain underwater-aware ability and model the complex distribution in real-world underwater scenes. INR Branch can learn robust feature representations from diverse underwater image, which reduces the difficulty of restoration for diffusion models.
2. Related Works
2.1. Underwater Image Enhancement
@@
Currently, existing UID methods can be briefly categorized into the physical and deep model-based approaches (Peng and Cosman, 2017; Peng et al., 2018; Tang et al., 2023; Peng et al., 2023; Li et al., 2020). Most UID methods based on the physical model utilize prior knowledge to establish models, such as water dark channel priors (Peng and Cosman, 2017), attenuation curve priors (Wang et al., 2018), fuzzy priors (Chiang and Chen, 2012). In addition, Akkaynak and Treibitz (Akkaynak and Treibitz, 2019b) proposed a method based on the revised physical imaging model. However, the depth map of the underwater scene is difficult to obtain. This leads to unstable performance of the this method, which usually suffers from severe color cast and artifacts. Therefore, the manually established priors restrain the model’s robustness and scalability under the complicated and varied circumstances. In recent years, underwater physical imaging models are gradually utilized in combination with data-driven methods (Li et al., 2021). Recently, deep learning-based methods (Tang et al., 2023; Peng et al., 2023; Li et al., 2020; Fu et al., 2022b) achieve acceptable performance, and some complex frameworks are proposed and achieve the-state-of-the-art performance (Peng and Cosman, 2017; Jr. et al., 2013; Zhao et al., 2023). The previous methods neglect the underwater imaging mechanism and rely only on the representation ability of deep networks. Ucolor (Li et al., 2021) combined the underwater physical imaging model in the raw space and designed a medium transmission guided model. ATDCnet (Mu et al., 2023a) attempted to use transmission maps to guide deep neural networks. Therefore, how to fully exploit the knowledge of physics in deep neural networks is a very crucial problem.
2.2. Diffusion Model
@@
Diffusion Probabilistic Models (DPMs) (Ho et al., 2020; Song et al., 2021) have been widely adopted for conditional image generation (Choi et al., 2021; Wang et al., 2023; Lu et al., 2023; Zhou et al., 2023, 2024). Saharia et al. (Saharia et al., 2022a) proposes Palette, which has demonstrated the excellent performance of diffusion models in the field of conditional image generation, including colorization, in-painting and JPEG restoration. Currently, diffusion models was employed to a variety of low-level visual tasks, and gained SOTA performance such as image restoration (Özdenizci and Legenstein, 2023), document enhancement(Yang et al., 2023b), and low light image enhancement (Yang et al., 2023c; Zhou et al., 2023). Recently, Tang et al. (Tang et al., 2023) presented an image enhancement approach with diffusion model in underwater scenes. WF-Diff (Zhao et al., 2024) proposed a underwater image enhancement framework based on frequency domain information and diffusion models. UIEDP propose a UIE framework with Diffusion Prior, treating UIE as a posterior distribution sampling process of clear images conditioned on degraded underwater inputs(Du et al., 2023). However, these methods fail to consider the physical properties and underwater imaging mechanisms in the diffusion process, limiting information completion capacity of diffusion models.
2.3. Implicit Neural Representation
Implicit Neural Representations (INRs) have emerged as an innovative technique garnering substantial attention for their ability to utilize coordinate-based Multi-Layer Perceptrons (MLPs) to model continuous-domain signals. Recently, some works have delved into the potential of INRs within the realm of two-dimensional image processing, including applications like image compression (Strümpler et al., 2022), and super-resolution(Chen et al., 2023). Chen et al. (Chen et al., 2021) first capitalized on INRs to achieve continuous image super-resolution. Subsequently, Lee et al. (Lee et al., 2022) proposed a dominant-frequency estimator that effectively predicts local texture information in natural images. Sun et al. (Sun et al., 2021) leveraged INRs to forecast continuous attributes based on the captured tomographic features. More recently, Quan et al. (Quan et al., 2023) introduced an INR-based inverse kernel prediction network to address the problem of single image defocus deblurring. Concurrently, Yang et al. (Yang et al., 2023a) exploited the controllable fitting capabilities of INRs to tackle the challenge of low-light image enhancement. Chen et al. (Chen et al., 2024) utilized INRs to learn common rain degradation representations from diverse rainy scenes. These efforts (Sun et al., 2021; Yang et al., 2023a; Chen et al., 2024) convincingly demonstrate the robust representation capabilities of INRs when applied to degraded images. Thus, inspired by the previous works, we explore the application of INRs to obtain robust feature representations for underwater image enhancement, aiming to reduce the difficulty of restoration for diffusion models.
Figure 2. The detailed architecture of our designed (a) physics-aware self-attention (PA-SA) and (b) gated multi-scale feed-forward network (GM-FFN).
3. Methodology
3.1. Overall Framework
@@
The overall framework of PA-Diff is shown in Fig. 1. PA-Diff is designed to leverage the physical properties of underwater imaging mechanisms to guide the diffusion process, which mainly consists of Physics Prior Generation (PPG) Branch, Implicit Neural Reconstruction (INR) Branch, and Physics-aware Diffusion Transformer (PDT) Branch. The PPG branch aims to generate the transmission map and the global background light as physics prior information by utilizing the modified koschmieder light scanning model. The transmission map is exploited as the confidence guidance for the PDT branch which enables our PA-Diff with underwater-aware ability. The global background light can be regarded as a prior condition on both the type and degree of degradation for the PDT Branch. Inspired by (Chen et al., 2024; Yang et al., 2023a) based on implicit neural representations that can encode an image as a continuous function and learn a more robust feature representation, we further introduce implicit neural representations to learn robust degradation representations from diverse underwater image, aiming to reduce the difficulty of restoration for diffusion models. The PDT branch employs the strong generation ability of diffusion models to restore underwater images, with the guidance of physical information. We adopt the diffusion process proposed in DDPM (Ho et al., 2020) to construct the PDT branch, which can be described as a forward diffusion process and a reverse diffusion process.
@@
(1) $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I),$
@@
(2) $x_{t}=\sqrt{\alpha_{t}}x_{t-1}+\sqrt{1-\alpha_{t}}\epsilon_{t-1},\quad\epsilon_{t-1}\sim\mathcal{N}(0,\mathcal{Z}).$
@@
(3) $q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha_{t}}}x_{0},(1-\bar{\alpha_{t}})I).$
@@
(4) $p_{\theta}(x_{t-1}|x_{t},x_{c})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},x_{c},t),\sigma_{t}^{2}\mathcal{Z}),$
where $x_{c}$ refers to the conditional image ${I}$ (underwater input image). $\boldsymbol{\mu}_{\theta}(x_{t},x_{c},t)$ and $\sigma_{t}^{2}$ are the mean and variance from the estimate of step t, respectively. We follow the setup of (Ho et al., 2020), they can be expressed as:
(5) $\mu_{\theta}(x_{t},x_{c},t)=\frac{1}{\sqrt{\alpha}_{t}}(x_{t}-\frac{\beta_{t}}{(1-\overline{\alpha}_{t})}\epsilon_{\theta}(x_{t},x_{c},t)),$
(6) $\sigma_{t}^{2}=\frac{1-\overline{\alpha}_{t-1}}{1-\overline{\alpha}_{t}}\beta_{t},$
@@
(7) $L_{dm}(\theta)=\|\epsilon-\epsilon_{\theta}(\sqrt{\overline{\alpha}_{t}}x_{0}+\sqrt{1-\overline{\alpha}_{t}}\epsilon,x_{c},t)\|.$
3.2. Physics Prior Generation
Mathematically, the underwater images can be expressed by the modified Koschmieder light scanning model (Akkaynak and Treibitz, 2019a):
(8) $I^{c}(x)=J^{c}(x)T^{c}(x)+(1-T^{c}(x))B^{c},$
where $c\in\{R,G,B\}$ is the color channel, $x$ indicates the spatial location of each pixel, $I^{c}$ is the observed underwater image, $J^{c}$ is the restored clean image, $T^{c}$ is the medium transmission map, and $B^{c}$ is the global background scattered light. In Physics Prior Generation (PPG) branch, our goal is to produce $T^{c}$ and $B^{c}$ from the underwater image $I^{c}$, exploiting the physics model.
Network architecture of PPG branch is shown in Fig. 1, and consists of two sub-networks: transmission map generation sub-network $\mathcal{T}(x)$ and background light generation sub-network $\mathcal{B}(x)$. The two sub-networks consist of duplicated Dynamic Convolutions (DC) (Chen et al., 2020). DC dynamically combines numerous parallel convolutional kernels according to their attention, thereby enhancing the model’s complexity and performance without augmenting the network’s depth or breadth. Given that the background light differs fundamentally from the transmission map, being independent of image content and instead encapsulating the global properties (Fu et al., 2022a) , for the purpose of generating the background light component, our background light generation sub-network initially employs a Gaussian blur block (GBB) to eliminate fine-grained content details and thereby extract the inherent global attributes.